OpenStudy (anonymous):

Expand f(x) = (x^3+2x^2+3x+4)^2 into polynomial form and use the power rule to find fprime(x)

OpenStudy (anonymous):

Polynomial means they want it in this kind of form x^3+2x^2+3x+4, notice this doesn't have the ( )^2 at the end. so like (x+y)^2=(x+y)*(x+y)=... Multiply the first term in the brackets,x in this case, by the first and second term in the brackets next door. this would give x*x and x*y. the same is done with y of the first brackets, y*x and y*y. do this same procedure for (x^3+2x^2+3x+4)*(x^3+2x^2+3x+4) very simple stuff just boring!!!!

OpenStudy (anonymous):

so foil basically?

OpenStudy (anonymous):

it'll take you 5 mins old sport!!¡¡ just multiplying stuff then the power rule.

OpenStudy (anonymous):

that's what i thought i just wasn't sure thank you

OpenStudy (anonymous):

ahh! hahaha thought by foil you meant you gave up :) my bad

OpenStudy (anonymous):

haha no way

OpenStudy (anonymous):

awesome